OpenStudy (anonymous):

Find the formula for the series: 2,1,0,1,2,1,0,1.....

OpenStudy (lgbasallote):

-1, -1, +1,+1,-1, -1, +1

OpenStudy (anonymous):

...directions were "find a formula for the nth term of the *sequence" my mistake

OpenStudy (lgbasallote):

hmmm..dunno how to put it in letters

OpenStudy (anonymous):

Any help would be appreciated. Even a shove in the right direction haha.

OpenStudy (anonymous):

f(3n+1) = 2, f(3n+2)=1, f(3n)=0

OpenStudy (anonymous):

wait, my answer is not right.

OpenStudy (anonymous):

f(4n+1) = 2, f(4n+2) = 1, f(4n+3)=0, f(4n)=1.

OpenStudy (anonymous):

You could say: f(n) = {2 if n mod 4 = 1, 1 if n mod 4 = 2, 0 if n mod 4 = 3, 1 if n mod 4 = 0}

OpenStudy (anonymous):

Ugh. I'm just not seeing it.

OpenStudy (anonymous):

How about this: \[f(n) = {1 \over2} (i (-i)^n-i i^n+2)\]

OpenStudy (anonymous):

Perfect! But, what does i stand for? :0

OpenStudy (anonymous):

well, if you don't have any imagination, there is probably a simpler answer.